# Bounds on decaying sterile neutrinos via magnetic dipole moment from COB intensity

### v2

@@ IV.3 Active-to-active neutrino magnetic moment
In the UV model, the active-to-active neutrino magnetic moment is naturally suppressed Beltrán et al. (2024). Active-sterile mixing can create an active-to-active neutrino magnetic moment in the broken phase parameters when the right-handed neutrinos are integrated out Beltrán et al. (2024). As such $d_{\nu\nu\gamma}\lesssim 2\times 10^{-8}\leavevmode\nobreak\ \text{\rm GeV}^{-1}$, we may anticipate that the active-to-active neutrino magnetic moments can satisfy the constraints from TEXONO Li et al. (2003), GEMMA Beda et al. (2010), LSND, and Borexino Agostini et al. (2017) in the inverse see-saw mechanism, where larger active-sterile mixing can occur. For the specified mass range $(0.12\leavevmode\nobreak\ \text{\rm eV}-0.7\leavevmode\nobreak\ \text{\rm eV})$ of active neutrinos, we are not able to obtain the specific intensity of photons due to the radiative decay of active neutrinos. As for this aforementioned mass range of active neutrinos, a very less number of photons can be generated, which cannot explain the COB anomalous intensity in the observed wavelength range. For that reason, we cannot observe any specific intensity of photons in the observed wavelength range $\lambda_{obs}=0.2\leavevmode\nobreak\ \mu m-1.2\leavevmode\nobreak\ \mu m$ because to explain the COB anomaly, we need minimum $4$ eV mass of sterile neutrinos. To summarize, we put constraints on the sterile-to-sterile and sterile-to-active transition magnetic dipole moment for ($1\,\rm keV-20\,\rm keV$) and ($4\,\rm eV-20\,\rm eV$) mass range of stile neutrinos, respectively by considering two COB measurements.
V Conclusion and Outlook
@@
Recent measurements of the cosmic optical background (COB) photon flux intensity by LORRI have revealed some excess intensity with no known sources. This excess may indicate a potential source of energetic photons beyond the Standard Model (BSM). In this study, we considered the possibility of decaying sterile neutrinos producing photons and explored the corresponding decay rates. Notably, we did not account for photon energy generation through cascades involving other decay products. We found the required decay rate of sterile neutrino and obtained upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments of the order $\mathcal{O}(10^{-22}-10^{-21})\leavevmode\nobreak\ \text{sec}^{-1}$ for the keV mass range of the decaying sterile neutrino and for the $\mathcal{O}$(eV) mass range of the decaying sterile neutrino, the decay rate should be in the range of $\mathcal{O}(10^{-23}-10^{-22})\leavevmode\nobreak\ \text{sec}^{-1}$. Here, we adopt a minimal framework involving sterile neutrinos using the effective field theory approach. The radiative decay of sterile neutrinos, resulting in photon production, is primarily governed by magnetic moments associated with sterile-to-sterile $(d_{NN\gamma})$ and sterile-to-active neutrino ($d_{\nu N\gamma}$) transitions. We calculate the $d_{NN\gamma}$ and $d_{\nu N\gamma}$ for various mass scales and decay rates. We find that, $d_{NN\gamma}$ take values of $3\times 10^{-11}\,\mu_{B}-10^{-9}\,\mu_{B}$ for sterile neutrinos of mass range $1\,\rm keV-40\,\rm keV$. Whereas sterile neutrinos, with mass range $4\,\rm eV-20\,\rm eV$, decaying via sterile-to-active transition magnetic moment $(d_{\nu\rm N\gamma})$ take values of $3.33\times 10^{-13}\,\mu_{B}-10^{-12}\,\mu_{B}$ for newly found COB measurement. In our work, we explain the bounds on sterile-to-sterile and sterile-to-active transition magnetic moments by considering the two COB measurements: one is $2.99\pm 2.03\,\rm nW/m^{2}/sr$ intensity, and another one is due to anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$. Here, we have plotted the specific intensity for both COB measurements. Motivated by this work, one can probe CUB, CXB, and CIB measurements to constrain the sterile-to-sterile and sterile-to-active transition magnetic dipole moment. We leave this as future work. Some upcoming experiments, such as GALEX and ULTRASAT satellites, can probe and constrain the decaying dark matter in the ultraviolet bandwidth Libanore and Kovetz (2025). Our minimal framework based on decaying sterile neutrinos of mass $\mathcal{O}$(keV) and $\mathcal{O}$(eV) can contribute to the unknown excess COB intensity, thus deriving upper bounds on the sterile-to-sterile and sterile-to-active transition magnetic moments.
VI Acknowledgement
@@
We sincerely acknowledge the usage of python packages scipy [^1] https://www.scipy.org/ and astropy [^2] https://www.astropy.org/index.html. We thank Rahul Kothari for his earlier collaboration on this project.

### v3

@@ IV.3 Active-to-active neutrino transition magnetic moment
In the UV model, the active-to-active neutrino transition magnetic moment is naturally suppressed Beltrán et al. (2024). Active-sterile mixing can create an active-to-active neutrino magnetic moment in the broken phase parameters when the right-handed neutrinos are integrated out Beltrán et al. (2024). As such $d_{\nu\nu\gamma}\lesssim 2\times 10^{-8}\leavevmode\nobreak\ \text{\rm GeV}^{-1}$, we may anticipate that the active-to-active neutrino magnetic moments can satisfy the constraints from TEXONO Li et al. (2003), GEMMA Beda et al. (2010), LSND, and Borexino Agostini et al. (2017) in the inverse see-saw mechanism, where larger active-sterile mixing can occur. For the specified mass range $(0.12\leavevmode\nobreak\ \text{\rm eV}-0.7\leavevmode\nobreak\ \text{\rm eV})$ of active neutrinos, we are not able to obtain the specific intensity of photons due to the radiative decay of active neutrinos. As for this aforementioned mass range of active neutrinos, a very less number of photons can be generated, which cannot explain the COB anomalous intensity in the observed wavelength range. For that reason, we cannot observe any specific intensity of photons in the observed wavelength range $\lambda_{obs}=0.2\leavevmode\nobreak\ \mu m-1.2\leavevmode\nobreak\ \mu m$ because to explain the COB anomaly, we need minimum $4$ eV mass of sterile neutrinos. Hence, the COB anomalous intensities can not be accounted for by the decay of active neutrinos into active neutrinos and photons.
V UV Origin of Small Mass Splittings in keV Sterile Neutrinos
Although our primary analysis adopts an effective field theory approach, a complete understanding of the small eV-scale mass splitting between keV sterile neutrinos benefits from a consistent ultraviolet (UV) completion. Several well-motivated frameworks exist that naturally explain the emergence of quasi-degenerate sterile neutrino masses without fine-tuning. Two particularly appealing UV approaches are as follows:
  1. 1.
    Froggatt–Nielsen Mechanism: In models with a $U(1)_{\rm FN}$ Froggatt–Nielsen symmetry Froggatt and Nielsen (1979); Pilaftsis and Underwood (2004), sterile neutrinos are assigned flavor-dependent charges under $U(1)_{\rm FN}$, and their Majorana masses arise from non-renormalizable operators suppressed by a high scale $\Lambda_{\rm FN}$ (the Froggatt–Nielsen (FN) cutoff scale, typically taken to be the GUT or Planck scale). The FN symmetry constrains the structure of the mass matrix, enforcing degeneracy at leading order. When the scalar field $S$ acquires a vacuum expectation value (VEV), the symmetry is spontaneously broken, generating mass terms suppressed by powers of the small parameter $\epsilon=\langle S\rangle/\Lambda_{\rm FN}$. Subleading contributions to the mass splitting may also arise from higher-order corrections or small perturbations consistent with the FN charge assignments. These effects induce naturally small mass splittings of the form
    $\Delta m\sim\epsilon^{n}\Lambda_{\rm FN}$
    where $n$ depends on the FN charge assignment and operator dimension.
    In particular, if two sterile neutrinos $N_{1}$ and $N_{2}$ are assigned suitable FN charges, they can acquire masses at the keV scale via higher-dimensional operators suppressed by $\epsilon=\langle S\rangle/M_{\rm GUT}$. The resulting Majorana mass terms exhibit the structure:
    $m_{N_{1,2}}\sim\epsilon M\sim\text{keV},\quad\Delta m_{12}\sim\epsilon^{2}M\sim\text{eV}$
    for $M\sim\text{TeV}$, where $M$ denotes the mass of a heavy mediator fermion $X$ that appears in the UV completion and is integrated out to generate the effective operator and $\epsilon\sim 10^{-6}$. This setup yields quasi-degenerate keV-scale sterile neutrinos with eV-level splittings—without requiring fine-tuning—making them viable dark matter candidates. The underlying process is diagrammatically shown in Figure 4.
[…]
    Figure 4: Majorana mass generation for sterile neutrino $N$ via two insertions of the FN scalar $S$, whose VEV breaks the $U(1)_{\rm FN}$ symmetry. The interaction is mediated by a heavy fermion $X$ of mass $M$.
  2. 2.
    Clockwork or Radiative Symmetry Breaking: In alternative frameworks, the degeneracy of sterile neutrino masses is protected by a discrete or continuous symmetry that is broken radiatively. This results in small loop-induced splittings between otherwise degenerate states. The clockwork mechanism, in particular, achieves exponential suppression of mass terms through a chain of fields with nearest-neighbor interactions Giudice and McCullough (2017); Maharana and Srivastava (2025); Hong et al. (2019); Hambye et al. (2004); Blanchet et al. (2010). When applied to sterile neutrinos, such setups can yield keV-scale mass eigenstates with eV-level splittings, making them phenomenologically viable and technically natural.
These UV completions demonstrate that small splittings among sterile neutrinos can emerge from structural features of the theory, rather than ad hoc fine-tuning, and are compatible with broader frameworks like seesaw and extra-dimensional models. Thus, the small $\mathcal{O}(\text{eV})$ mass difference between keV sterile states—essential for oscillation and decay phenomenology—can be realized in a variety of UV-consistent ways.
VI Conclusion and Outlook
@@
Recent measurements of the cosmic optical background (COB) photon flux intensity by LORRI have revealed some excess intensity with no known sources. This excess may indicate a potential source of energetic photons beyond the Standard Model (BSM). In this study, we investigated the possibility of sterile neutrinos decaying into photons and explored the corresponding decay widths. Notably, we did not account for photon energy generation through cascades involving other decay products. We found the required decay width of the sterile neutrino and obtained upper bounds on the sterile-to-sterile transition magnetic moments of the order $\mathcal{O}(10^{-22}-10^{-21})\leavevmode\nobreak\ \text{sec}^{-1}$ for the keV mass range of the decaying sterile neutrino. Here, we adopt a minimal framework involving sterile neutrinos using the effective field theory approach. The radiative decay of sterile neutrinos, resulting in photon production, is primarily governed by magnetic moments associated with the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$. We calculate the $d_{NN\gamma}$ for the different $\Delta m$ values, which vary from $2-10\,\rm eV$. We find that, $d_{NN\gamma}$ take values of $3\times 10^{-12}\,\rm eV^{-1}-10^{-9}\,\rm eV^{-1}$ for sterile neutrinos of mass range $1\,\rm keV-40\,\rm keV$. In our work, we explain the bounds on sterile-to-sterile transition magnetic moments by considering the two COB measurements: one is $2.99\pm 2.03\,\rm nW/m^{2}/sr$ intensity, and the other one is due to anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$. Here, we have plotted the specific intensity for both COB measurements. Motivated by this work, one can probe CUB, CXB, and CIB measurements to constrain the sterile-to-sterile transition magnetic moment. We leave this as future work. Some upcoming experiments, such as GALEX and ULTRASAT satellites, can probe and constrain the decaying dark matter in the ultraviolet bandwidth Libanore and Kovetz (2025). Our minimal framework based on decaying sterile neutrinos of mass $\mathcal{O}$(keV) can contribute to the unknown excess COB intensity, thus deriving upper bounds on the sterile-to-sterile transition magnetic moments. In conclusion, our analysis establishes that the anomalous rest-frame intensity of the Cosmic Optical Background enables constraints on the sterile-to-sterile transition magnetic moment, $d_{NN\gamma}$, while scenarios involving sterile-to-active or active-to-active transitions remain disfavored within this framework.
VII Acknowledgement
@@
We sincerely acknowledge the usage of python packages scipy [^1] https://www.scipy.org/ and astropy [^2] https://www.astropy.org/index.html. We thank Rahul Kothari for his earlier collaboration on this project. H. H. acknowledges support from the Ministry of Higher Education, Government of India. T. S. is supported by the National Natural Science Foundation of China under Grants No. 12475094, No. 12135006, and No. 12075097.